OpenStudy (anonymous):

Help Please LMNO is a parallelogram. If NM=x+15 and OL=3x+5, find the values of x and then find NM and OL a.)x=7;NM=20;OL=22 b.)x=5;NM=20;OL=20 c.)x=7;NM=22;OL=22 d.)x=5;NM=22;OL=20

OpenStudy (anonymous):



OpenStudy (nurali):

If NM=OL Then x + 15 = 3x + 5 Solve for x x = 5

OpenStudy (nurali):

NM=x+15 NM=5+15 NM=20

OpenStudy (anonymous):

So it B thank you can you maybe help me on some more questions??

OpenStudy (nurali):

OL=3x+5 OL=3(5)+5 OL=20

OpenStudy (nurali):

Anytime.